OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1). − −

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

the options are -7/3 , 3/7 , 7/3 , and 3/7

OpenStudy (anonymous):

are we assuming a liner line here???

OpenStudy (anonymous):

idk it says Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1). , and those are the opption they gave me

OpenStudy (anonymous):

ok am assuming a liner line here. so y=mx+b m=change in y / change in x

OpenStudy (anonymous):

so is it the 3rd opptionn?

OpenStudy (anonymous):

sorry 7/3

OpenStudy (anonymous):

so -6 - 1)/ -6-(-3) =7/3